OpenStudy (anonymous):

The following table shows the values of f(x) and g(x) for different values of x: x 1, 2, 3, 4, 5 f(x) 4, 16, 64, 256, 1024 g(x) -1, -1, -1, -1, -1 Which of the following best describes the function f(x) + g(x)? A. 4x - 1x B. -1(4x) C. 4x - 1 D. 4x - 1 I really need help on this one.

OpenStudy (anonymous):

Well let's go ahead and draw some graphs. We can use the points (x, f(x))'s to draw the coordinates to illustrate the graph of (f) We can use the points (x, g(x))'s to draw the coordinates to illustrate the graph of (g)

OpenStudy (anonymous):

And since the numbers are so high, graphing it will be a little tricky. I just wanted you to see how it looked when you graph two functions and then add them. Do you know the basics of adding functions?

OpenStudy (anonymous):

Sort of. Can you go through the basics a little?

OpenStudy (anonymous):

I can't seem to figure out a function that does (1, 4), (2, 16), (3, 64)...

OpenStudy (anonymous):

We need to figure out the functions first before we can add them together.

OpenStudy (anonymous):

Gimme a few minutes.

OpenStudy (anonymous):

I'm going to figure this out, no worries =0)

OpenStudy (anonymous):

No rush man!

OpenStudy (anonymous):

Alright so our first function is - since we know the values of x are 1, 2, 3, 4, etc. and the values of f(x) are 4, 16, 64, 256; that our f equation is:\[f(x) = 4^{x}\] Does that make sense so far?

OpenStudy (anonymous):

That's not a negative in the first sentence lol

OpenStudy (anonymous):

Yes, except the dash in the first sentence. Is that just a blank? :P

OpenStudy (anonymous):

Don't worry about the dash, that's the way I separate ideas in sentences lol, I shouldn't do that in Math haha

OpenStudy (anonymous):

Alright so our first function, since we know the values of x are 1, 2, 3, 4, etc. and the values of f(x) are 4, 16, 64, 256; that our f equation is:

OpenStudy (anonymous):

Read it like that

OpenStudy (anonymous):

Yeah, I understand.

OpenStudy (anonymous):

Sorry about that. So our first function is f(x) = x^4 And our g(x) function has the points (1, -1), (2, -1), (3, -1), etc. Which can be shown by the function g(x) = -1

OpenStudy (anonymous):

\[f(x) = 4^{x}\]\[g(x) = -1\] So we need to know what \[f(x) + g(x) \] is right?

OpenStudy (anonymous):

What are your answer choices again? Choice A: 4x - 1x doesn't make much sense because this is just 3x, and usually answers/solutions will never put a 1 in front of a solitary x...

OpenStudy (anonymous):

Are you sure their not 4^x - 1^x and -1(4^x) etc...?

OpenStudy (anonymous):

And you have the same answer for C and D

OpenStudy (anonymous):

C. 4x - 1 D. 4x - 1 ....

OpenStudy (anonymous):

Oh, no. C has 1 as an exponent.

OpenStudy (anonymous):

So it would be more like 4x - ^1

OpenStudy (anonymous):

Sorry.

OpenStudy (anonymous):

4x - ^1 makes absolutely no sense whatsoever.

OpenStudy (anonymous):

Like this? \[4x ^{-1}\]

OpenStudy (anonymous):

Yeah, But that only leaves B then.

OpenStudy (anonymous):

And B seems the most logical answer anyways.

OpenStudy (anonymous):

-1(4x) is just equal to -4x. And that is not what 4 raised to the x power + (-1) would be at all.

OpenStudy (anonymous):

You must be missing an exponent, or some sign in there. Is it not\[-1(4^{x})\]

OpenStudy (anonymous):

Because that is Completely different than -1(4x)

OpenStudy (anonymous):

... Crap, you're right. :| I'm sorry, I didn't notice the exponents. on C and B. Let me re-write them... A. 4^x - 1x B. -1(4^x) C. 4^x - ^1 D. 4^x - 1

OpenStudy (anonymous):

Anyways f(x) + g(x) = (f + g)(x) \[4^{x}+(-1) = 4^{x}-1\]

OpenStudy (anonymous):

Thank you. Medal for you!